joy was a thrice as old as jill five years ago. Now, she is twice as old as jill. How old are they now ?
Answer by lwsshak3(11628)

joy was a thrice as old as jill five years ago. Now, she is twice as old as jill. How old are they now ?
***
now:
x=Jill"s age
2x=Joy's age
..
5 yrs ago:
x-5=Jill"s age
2x-5=Joy's age
2x-5=3(x-5)
2x-5=3x-15
x=10
2x=20
..
Jill"s age now=10
Joy's age now=20